What is 8x - 5y = 15 in slope intercept form
Answer by nerdybill(7384)

What is 8x - 5y = 15 in slope intercept form
.
"Slope-intercept" form is:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
8x - 5y = 15
-5y = -8x + 15
5y = 8x - 15
y = (8/5)x - 3 (this is what they're looking for)